OpenStudy (anonymous):

(sinx-cosx)^(2)- (sinx+cosx)^(2) PLEASE HELP!

OpenStudy (solomonzelman):

has some tips and formulas in it.

OpenStudy (anonymous):

(sin x - cos x)^2 = sin^2(x) - 2 sinx cosx + cos^2(x) (sin x + cos x)^2 = sin^2 (x) + 2sinx cox + cos^2(x) so the first expression minues the second expression is just: - 4sinx cos x

OpenStudy (anonymous):

You could have also done this problem by realizing that sin^2x + cos^2x = 1 so the first expression is just 1-2sin x cos x And the same for the second expression, you get 1+2sinx cos x Now subtract, and you'll still get -4 sin x cos x

OpenStudy (anonymous):

Your welcome.

OpenStudy (raden):

alternative : use the difference square formula : m^2 - n^2 = (m+n)(m-n) now, (sinx-cosx)^(2)- (sinx+cosx)^(2) = (sinx-cosx+sinx+cosx)(sinx-cosx-(sinx+cosx)) = 2sinx(-2cosx) =-4sinxcos

OpenStudy (raden):

-4sinxcosx

OpenStudy (anonymous):

THANK YOU ALL SO MUCH SERIOUSLY LIFESAVERS....